OpenStudy (anonymous):

The following is an arithmetic sequence 1/2, 1/ 3/2, 2 true or false. I wanna say false for this one? The following is an arithmetic sequence. -3, 3, -3, 3 False True and I wanna say false for this one am i correct?

jimthompson5910 (jim_thompson5910):

Seems like there's a typo in " 1/2, 1/ 3/2, 2"

OpenStudy (anonymous):

1/2, 1/ 3/2, 2 or 1/2, 1, 3/2, 2

OpenStudy (anonymous):

see 1/2 +1/2 =1 1+1/2 = 3/2 3/2 +1/2 = 2 thus the common difference is 1/2 and constant so an arithmetic sequence..similarly can you check the second sequence you posted

OpenStudy (cp9454):

in A.P the comman difference is same for all the terms. i.e \[a _{n}-a _{n-1} = d\] always , you can check by putting n=1,2,3,4 .......n any number what so ever. but remember here n represents the number of terms of series. and in your case the comman difference is 1/2.

OpenStudy (anonymous):

it's supposed to be 1/2, 1, 3/2, 2

OpenStudy (tkhunny):

And? Subtract each successive pair and see what you get? 1 - 1/2 = ?? 3/2 - 1 = ?? 2 - 3/2 = ??

OpenStudy (anonymous):

.5

OpenStudy (anonymous):

so yes okie thanks!